The population in Moldova is 5 mil. There were 190,000 deaths during the year 2018. There were 10,000 deaths from cancer and 60,000 deaths from cvd. There were 80,000 people with cancer. Show all work and include all units.

  1. What was the annual mortality rate in Moldova in 2018 per 1,000? Annual mortality rate from Cancer in Moldova in 2018 and the annual mortality rate from cvd in 2018? Proportionate mortality from cancer in 2018?
Answer #1

(1)

Question:

What was the annual mortality rate in Moldova in 2018 per 1,000?

Answer:

The annual mortality rate in Moldova in 2018 per 1,000 = (Number of deaths/ Number of people in the population) X 1000

= (190,000/5,000,000) X 1000

= 38 deaths per 1,000 unit of population in 2018

(2)

Question:

Annual mortality rate from Cancer in Moldova in 2018

Answer:

Annual mortality rate from Cancer in Moldova in 2018 = Number of deaths from Cancer/ Number of people in the population

= 10,0000/ 5,000,000

= 0.002 deaths per 1 unit of population in 2018

(3)

Question:

the annual mortality rate from cvd in 2018

Answer:

the annual mortality rate from cvd in 2018 = Number of deaths from cvd/ Number of people in the population

= 60,000/5,000,000

= 0.012 deaths per 1 unit of population in 2018

(4)

Question:

Proportionate mortality from cancer in 2018?

Answer:

Proportionate mortality from cancer in 2018 = Number of deaths from cancer/ Total number of deaths

= 10,000/190,000

= 0.0526 deaths from cancer per 1 unit of total death in 2018